OpenStudy (anonymous):

HELP!!! 2x-y=-7 5x+y=-14 Find x and y??

OpenStudy (anonymous):

If I did this correctly, not quite sure it is thisl 2x-y=-7 5x+y=-14 7x=-7 x=-1 Could anyone confirm this?

OpenStudy (anonymous):

i got x is equal to positive one. ill double check on paper.

OpenStudy (anonymous):

yeah that looks right but there is another step?

OpenStudy (anonymous):

then put your answer for x into one of the original equations and solve for y

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

Well @xkaykayx92 7x=-7 /7 /7 x = -1 negative divided by a positive equals a negative, and vise versa.

OpenStudy (anonymous):

Well @dread907 i got 7x=7 not negative 7 so yeah.

OpenStudy (anonymous):

-2-y=-7 whats the next step?

OpenStudy (anonymous):

thats after i did 2(-1)-y=-7

OpenStudy (anonymous):

How did you manage that? Unless you did it opposite of me because I went -14 - (-)7 = 14 + 7 = -7 Did you do -7 - (-14) = -7 + 14 = 7 ?

OpenStudy (anonymous):

Combine like terms. so; y = -5

OpenStudy (anonymous):

-2 and -7 being the like terms.

OpenStudy (anonymous):

thanks! You guys are so amazing :) @dread907 @xkaykayx92

OpenStudy (anonymous):

After that plug x and y into the other equation.

OpenStudy (anonymous):

okay i have another problem ahah

OpenStudy (anonymous):

whats the problem @katie__michelle

OpenStudy (anonymous):

4x+y=6 -2x+4y=6

OpenStudy (anonymous):

i got x=1 and y=2 @dread907 do you agree with that?

OpenStudy (anonymous):

Yes, but it would help Katie if you showed all of your work instead of simply giving an answer.

OpenStudy (anonymous):

@dread907 thanks haha

OpenStudy (anonymous):

okay take 4x+y=6 and make it y=-4x+6 then plug that equation into the other one to solve for x -2x+4(-4x+6)=6 and you get -2x-16x+24=6 then solve for x -18x=-18 x=1 then plug x into an orginal equation to solve for y 4x+y=6 4(1)+y=6 y=2

OpenStudy (anonymous):

Good job on the explanation! Don't forget to close your question if your done Katie.

OpenStudy (anonymous):

okay thanks guys